amengz2:

the displacement of a mass suspended on a spring is modeled by the function y=5sin(pi t), where y is measured in inches and t in seconds. the amplitude, period, and frequency of the emotion of the mass are given by .

Mercury:

Answer choices: a. amplitude=2, period = 2 and frequency = 2 b. amplitude = 5, period = 2, and frequency = 1/2 c. amplitude = 5, period = 1 and frequency = 1 d. amplitude = 5, period = 1/2 and frequency = 2 compare to the general formula y = a(b(x-h)) + k, where |a| is the amplitude, b = 2pi/(period), and frequency = 1/period. h and k are horizontal and vertical shifts, but we don't have them here. find your a value (amplitude) find your b-value, and plug into the formula b = 2pi/period to find period then calculate 1/period to find frequency